OpenStudy (bananas1234):

Find the quotient. 14x^3 - 21x ------------- 7x

OpenStudy (michele_laino):

we can factor out 7x at the numerator, so we get: \[\Large \frac{{14{x^3} - 21x}}{{7x}} = \frac{{7x\left( {2{x^2} - 3} \right)}}{{7x}} = ...?\]

OpenStudy (bananas1234):

14^3 -3? @Michele_Laino